### v1

# AnyCap Project: A Unified Framework, Dataset, and Benchmark for Controllable Omni-modal Captioning

Abstract
Controllable captioning is essential for precise multimodal alignment and instruction following, yet existing models often lack fine-grained control and reliable evaluation protocols. To comprehensively address this gap, we introduce the AnyCap Project, developing an integrated solution spanning model, dataset, and evaluation. Specifically, we propose AnyCapModel (ACM), a lightweight plug-and-play framework that enhances the controllability of existing foundation models for omni-modal captioning, avoiding the training of base model. ACM leverages the original caption from base models, while incorporating user instructions and modality features to produce improved captions. In addition, to make up for the data deficiencies of controllable multimodal captioning, we construct AnyCapDataset (ACD), which covers $3$ modalities, $28$ types of user instructions, and $300$k high-quality data entries. We further propose AnyCapEval, a novel benchmark that offers more reliable evaluation metrics for controllable captioning by decoupling content accuracy and stylistic fidelity. ACM significantly improves caption quality across a diverse set of base models on AnyCapEval. Notably, ACM-8B boosts GPT-4o’s content scores by $45\%$ and style scores by $12\%$. It also shows substantial gains on widely-used benchmarks such as MIA-Bench and VidCapBench.
1 Introduction
Figure 1: The overall framework of ACM. Given image, video, and audio inputs, base captioning models first generate initial outputs. ACM refines these outputs by aligning modality features with specific user instructions, enabling controllable captioning without retraining the base models.
Multimodal large language models (MLLMs) are rapidly evolving towards omni-modal intelligence, processing diverse modalities including images, videos, and audio. Textual captions serve as a universal interface linking these modalities, benefiting numerous tasks such as retrieval, question answering, and content generation [1, 2, 3, 4, 5, 6]. Recently, controllability in captioning has garnered increasing attention. It refers to the ability to generate captions that precisely follow user instructions, such as emphasizing specific aspects or adopting designated stylistic expressions. Such control is crucial for tailoring captions to user needs and enhancing downstream tasks [7, 8, 9]. Yet, achieving precise and reliable omni-modal controllable captioning remains a significant challenge, largely constrained by limitations in current framework, data, and evaluation approaches [10, 11, 12, 13, 14].
Controllable captioning across diverse modalities faces three main challenges: (1) Limited controllability capabilities of open-source models. Existing caption models typically rely on rigid control signals, such as soft prompts [13] or bounding boxes [15, 14], to adjust the fine-granularity of the caption. While this allows certain control, it limits flexibility and reduces the variety of generated results. Natural language prompts provide a more flexible way to guide captions, yet current methods often perform well only on specific aspects (e.g., detailed descriptions) due to limited and narrow training data. Retraining these models to support broader control is expensive and may weaken their overall language ability. (2) Lack of controllable omni-modal caption datasets. Manual annotation for controllable captioning is costly and labor-intensive. Meanwhile, due to the lack of low-cost and quality-assured data generation pipelines, generating large-scale data using models remains challenging. (3) Lack of appropriate evaluation benchmarks and metrics. Classic metrics (e.g., BLEU [16], CIDEr [17]) ignore content accuracy and stylistic adherence, whereas LLM-based scorers suffer from high variance, stylistic bias, and poor diagnostic power [18, 19, 20, 21, 22, 23, 24, 25].
To address these challenges, we propose AnyCapModel (ACM), a plug-and-play framework that enhances fine-grained controllability in omni-modal caption generation with frozen base models. By aligning user instructions, modality features, and initial captions, ACM produces improved captions to meet diverse control demands (Fig. 1). Without expensive training of base models, it offers a general and low-cost path toward controlled multimodal captioning.
For the lack of training data for controllable captioning, we introduce AnyCapDataset (ACD), a large-scale omni-modal dataset spanning images, videos, and audio across 28 distinct control dimensions. Each sample includes customized control signals and a pair of captions, one preferred and one rejected, clearly reflecting differences in control quality. Crucially, this preference-based design only requires that chosen caption outperform their rejected one, largely lowering annotation difficulty. This design supports learning precise alignment with diverse user preferences, which is directly compatible with our training paradigm and applicable to other preference-based methods [26, 27, 28, 29, 30, 31].
To enable a more reliable assessment for controllable captioning, we design AnyCapEval, a benchmark that assesses content and style dimensions separately. We propose Keypoint Density (KPD) and detailed scoring criteria, providing more objective and precise evaluation metrics compared to existing methods. Specifically, the benchmark evaluates both Content that focuses on adherence to control instructions, and Style that captures expressive fidelity and hallucination reduction. This two-dimensional evaluation gives rise to improved accuracy and reduced variance in assessment.
Extensive experiments demonstrate that ACM significantly improves controllability across image, video, and audio modalities, benefiting both proprietary models and various open-source alternatives. Specifically, integrating ACM-2B with GPT-4o [1] improves content controllability by 37% and style controllability by 8%. Moreover, the larger ACM-8B empowers certain open-source models, such as InternVL2.5-8B [32], to match or even surpass proprietary GPT-4o across multiple control dimensions. It consistently shows improvements on widely-used benchmarks including MIA-Bench [22] and VidCapBench [23]. In summary, our main contributions are:
(1) We introduce ACM, a versatile plug-and-play framework that boosts caption controllability for frozen base models across image, video, and audio modalities. Given the scarcity of controllable caption data, we construct ACD, a large-scale omni-modal dataset with preference pairs tailored for controllable captioning, to the best of our knowledge.
(2) We propose AnyCapEval, a novel evaluation benchmark designed for controllable captioning. It enhances assessment reliability by providing both content and style related metrics.
(3) We provide extensive experiments to show that ACM greatly improves controllability and reduces hallucinations across diverse modalities and models, achieving significant gains on both AnyCapEval and the widely-used benchmarks.
2 Related Work
Controllable captioning. Controllable captioning conditions text generation on user-instructions to dictate content and style. Early efforts in this field primarily focused on incorporating predefined and often low-level control signals. For example, researchers explored controlling caption generation based on explicit properties of the output text, such as desired caption length [11], or by conditioning on specific visual information from the image [33, 34]. This included leveraging particular visual regions or objects, sometimes specified via bounding boxes [7], or incorporating specific aspects for video captioning [35, 36]. These methods, while pioneering, lacked the expressiveness and intuitiveness of natural-language instructions that align closely with human preferences. More recent efforts leveraging MLLMs have significantly advanced image captioning, yet still often produce global and modality-specific captions, limiting their applicability to omni-modal settings [8, 9, 12, 37, 38, 39, 40, 13, 41, 42, 43, 44]. In contrast, our ACM supports flexible, language-based controls across images, videos, and audio, which can be easily extended to additional modalities.
Captioning benchmarks. Existing captioning benchmarks are inadequate for precisely evaluating caption controllability. Metrics originally developed for machine translation, such as BLEU [16] and CIDEr [17], poorly reflect semantic accuracy and controllability. Critically, their reliance on n-gram overlap fails to verify adherence to specific control signals, focusing instead on general similarity. While evaluations leveraging large language models (LLMs) can offer some insight into semantic coherence [18, 22, 45, 46, 47, 48, 49, 50, 51], they often tend to be unreliable due to significant scoring variability and potential biases [52, 53]. Similarly, evaluating adherence to fine-grained control using LLMs is often subjective and inconsistent. To address these limitations, we introduce a more fine-grained benchmark that employs keypoint density (KPD) and detailed scoring criteria, rigorously assessing controllability from both content and style dimensions.
3 AnyCap Project
In Sec. 3.1, we first propose AnyCapModel (ACM), a plug-and-play approach to help existing foundation models with caption controllability without fine-tuning. In Sec. 3.2, we present AnyCapDataset (ACD), a comprehensive dataset featuring a rich diversity of controllable captions. Finally, in Sec. 3.3, we detail the AnyCapEval benchmark for caption evaluation of control accuracy and hallucination.
3.1 AnyCapModel
Figure 2: The workflow of AnyCapModel. (a) Integrating ACM into frozen base models enhances caption controllability across multiple modalities. (b) ACM aligns modality features, initial captions, and user instructions to produce refined, instruction-compliant captions.
Workflow. As illustrated in Fig. 2(a), given a modality input $m$ from possible modalities such as image ($m^{\text{img}}$), video ($m^{\text{vid}}$), or audio ($m^{\text{aud}}$), alongside a user-specified instruction $q$, an initial caption $y_{0}$ is first generated by a frozen base model $\mathcal{M}_{g}$. Subsequently, our alignment framework ACM ($\mathcal{M}_{a}$) refines this caption:
$y_{c}=\mathcal{M}_{a}(m,q,y_{0}).$ (1)
As shown in Fig. 2(b), modality-specific encoders (e.g., InternViT [32] for $m^{\text{img}}$ and $m^{\text{vid}}$, and EAT [54] for $m^{\text{aud}}$) extract feature embeddings from respective modality inputs. These modality embeddings are projected into a shared semantic space via modality-specific linear transformations (MLPs). Concurrently, $q$ and $y_{0}$ are tokenized and embedded into textual embeddings. Finally, modality and textual embeddings are concatenated and input into an autoregressive language model to produce a refined, instruction-compliant caption $y_{c}$.
Training strategy. We adopt the residual correction training strategy [55]. This approach focuses on refining existing captions rather than generating perfectly accurate ones from scratch. To effectively train the model, ACM learns how to correct uncontrolled or hallucinated outputs by contrasting them with controlled, accurate examples. During training, we include approximately $40\%$ of data samples where initial captions already match desired controlled captions or are factually correct. This inclusion explicitly guides the model in recognizing instances requiring minimal or no correction, thereby enhancing its overall alignment capability.
Training setting. We employ the AdamW optimizer for $3$ epochs with a learning rate of $1\times 10^{-6}$, a cosine learning rate schedule ($0.03$ warmup ratio), weight decay of $0.01$, and a global batch size of $256$, using bfloat16 mixed-precision. During ACM training, the external modality backbones are frozen, while the ACM’s internal components including its language model are updated.
3.2 AnyCapDataset
Definition. This dataset is constructed to train omni-modal systems that generate descriptions under explicit user instruction. ACD adopts a carefully designed triplet structure:
$D_{\text{ACD{}}}:=\left\{(q_{i},c_{i},a_{i})\right\}_{i=1}^{N},$ (2)
where $q_{i}$ denotes an instruction that specifies the required properties of a caption. For each instance, $c_{i}$ is a high-quality caption adhering to the instruction, while $a_{i}$ is a suboptimal caption that may exhibit minor deficiencies in factual accuracy, level of detail, or compliance with the instruction.
Table 1: Supported instruction types across image, video, and audio modalities in AnyCapDataset. Checkmarks (✓) indicate where each control is applicable. Content controls regulate what is said, whereas Style controls regulate how it is articulated.
Instruction Type Description Image Video Audio
Content
Background (Bkg) Provide or suppress scene–background details
Event (Evt) Require mention of an event or temporal change
Instance (Ins) Emphasise or ignore specific entities; enable inter-entity comparison
Instance Action (IAct) Describe the motion state of a designated instance
Instance Appearance (IApp) Characterise the visual attributes of a designated instance
Instance Position (IPos) Specify spatial position of an instance within the scene
Movement (Mov) Specify camera motion type (e.g., pan, zoom, dolly)
Perspective (Per) Require a particular viewpoint of an object/person
Region (Reg) Restrict description to a specified image/video region
Style
Brief (Brf) Produce a concise rendition with minimal elaboration
Detail (Det) Regulate the required level of descriptive granularity
Genre (Gen) Adopt a literary form (e.g., poem (Poe), narrative (Nar))
Length (Len) Constrain caption size (words or sentences)
Theme (Thm) Conform to a designated linguistic style
Instruction types. The instruction types (shown in Table 1) are systematically identified by surveying caption-related literature to find common control categories and analyzing typical requirements of downstream applications. This systematic process allows us to curate a set of instruction types aiming for balanced diversity and practicality, facilitating precise control over caption generation.
Construction pipeline. For each multimodal sample, a user instruction $q$ and a related compliant caption $c$ are generated in a single step. For each instruction type, we design a detailed prompt including: (1) task description, (2) multimodal input, (3) original reference title if available, and (4) several high-quality examples that illustrate the desired content or style constraints. These prompts guide the model to produce well-aligned $(q,c)$ pairs. Further details can be found in the Appendix A.
We primarily use open-source models like InternVL2.5 [32], and apply proprietary models only for more complex cases such as perspective, balancing quality and cost. Before large-scale generation, each prompt goes through a strict validation phase: we sample about $20$ instances per modality and instruction type, and manually verify that the captions strictly follow the user instruction, contain no hallucinations, meet formatting requirements, and preserve the input’s unique features. Only prompts with a $100\%$ pass rate are used for batch generation. To generate suboptimal captions $a$, we either remove guidance or add a controlled degradation instruction (e.g., missing details, factual errors, or style violations) based on the given $(q,c)$ pair. Additionally, after data generation, a random $5\%$ sample underwent manual review, finding over $95\%$ compliance with human preferences.
Significant human and temporal resources are dedicated to establishing a scalable and cost-effective pipeline. This investment has yielded a robust dataset foundation, substantially reducing costs for subsequent research and enabling the low-cost, large-scale generation of high-quality caption data.
Statistics. Following the construction pipeline described above, ACD ultimately comprises approximately $300$k $(q,c,a)$ triplets, derived and augmented from around $75$k original samples. These original samples are collected from multiple public datasets [56, 57, 58, 59, 60, 61]. ACD spans three modalities: image ($125$k samples), video ($100$k samples), and audio ($75$k samples). The average length of optimal captions $c$ is $62$ words, while suboptimal captions $a$ average $57$ words. Further details can be found in the Appendix A.
3.3 AnyCapEval
Figure 3: Evaluation methodology of AnyCapEval. (a) Examples demonstrating content evaluation via Keypoint Density (KPD) and style scoring rules. (b) KPD correlation analysis showing superior reliability. (c) Radar chart depicting performance gains from ACM integration.
We introduce AnyCapEval, a benchmark designed for systematic evaluation of controllability in omni-modal caption generation. Control requirements include two types. (1) Content assesses whether captions accurately follow control conditions and maintain semantic relevance. (2) Style evaluates consistency with human references in narrative manner, structure, and fluency.
Content evaluation. Given a reference caption $r$ and a user instruction $q$, we first annotate a key-point set $\mathcal{K}_{r,q}=\{k_{1},k_{2},\dots,k_{n}\}$ that exhaustively covers the information required by $q$. The candidate caption is denoted as $\hat{y}$. An automatic matcher based on GPT-4o [1] identifies the subset of key points present in $\hat{y}$ (Fig. 3(a) presents illustrative examples). To normalise for verbosity, we introduce the key-point density (KPD):
$\text{KPD}(\hat{y};{q})\;=\;\frac{N_{\text{match}}}{L_{\text{words}}(\hat{y})}\times 100,$ (3)
where $N_{\text{match}}$ denotes the count of key points from $\mathcal{K}_{r,q}$ present in $\hat{y}$, and $L_{\text{words}}(\hat{y})$is the total number of words in $\hat{y}$. KPD quantifies the effective information rate—matched key points per $100$ words.
While a naive alternative might simply count all extracted information points, KPD offers a more robust measure of effective information rate. Empirically, the Pearson correlation between this raw count ($R_{\text{all}}$) and human-judged relevance is weak, whereas the correlation for KPD in Eq. 3 is much stronger, as shown in Fig. 3(b). Length alone ($L_{\text{words}}$) already captures noise, so density provides a more stable, scale-invariant signal. Further details can be found in the Appendix B.
Style evaluation. For style evaluation, GPT-4o compares the candidate caption $\hat{y}$ with the reference $r$ under the user instruction $q$ and assigns a discrete score $s(\hat{y})\in\{0,\dots,4\}$. The scoring criteria are carefully defined as follows: (0) severely deviates from $q$ or largely hallucinated/incorrect; (1) deviates from $q$ or contains many hallucinations; (2) slightly deviates from $q$ or contains some hallucination; (3) highly similar to $r$ and hallucination-free; (4) outperforms $r$ while remaining hallucination-free. (see Fig. 3(a) for examples).
This fine-grained scoring approach, guided by a detailed, structured prompt, supplies GPT-4o with the candidate $\hat{y}$, reference $r$, user instruction $q$, and explicit criteria along three axes—semantic similarity, hallucination severity, and stylistic conformity—for each level $0\to 4$. This tight rubric sharply reduces evaluation variance and subjective bias. Further details can be found in the Appendix B.
4 Experiments
4.1 Experimental Setup
We train the AnyCapModel (ACM) framework in two variants (2B and 8B) on our proposed AnyCapDataset (ACD), which comprises approximately 125K images, 100K videos, and 50K audio samples. Training is conducted using 32 NVIDIA A100 (80GB) GPUs, requiring just 6 hours for the 2B model and 21 hours for the 8B model. More training details are provided in the Appendix C.
4.2 Evaluation on Controllable Caption
Table 2: Results of controllable image captioning on AnyCapEval. Abbreviations like IPos. are different instruction types in Table 1. Content and style values are computed using KPD and style scores from Sec. 3.3, respectively. ACM significantly improves content accuracy and stylistic fidelity.
Model Content Style
IPos. IApp. Ins. Per. Avg. Brf. Det. Thm. Poe. Nar. Avg.
Proprietary Models
GPT-4o 1.55 3.25 3.84 2.94 2.89 1.91 1.86 2.59 2.74 2.59 2.26
 +ACM-2B 3.01 (+1.46) 4.32 (+1.07) 4.55 (+0.71) 4.57 (+1.62) 4.11 (+1.22) 2.29 (+0.38) 1.98 (+0.12) 2.38 (-0.21) 3.02 (+0.29) 2.78 (+0.20) 2.46 (+0.19)
 +ACM-8B 3.48 (+1.93) 4.81 (+1.56) 4.89 (+1.06) 5.00 (+2.05) 4.54 (+1.65) 2.38 (+0.47) 2.35 (+0.49) 2.82 (+0.24) 3.10 (+0.37) 2.87 (+0.28) 2.65 (+0.39)
Open-source Models
InternVL2.5-8B 1.51 3.43 4.54 2.68 3.04 2.13 1.92 1.94 2.08 2.52 2.12
 +ACM-2B 3.22 (+1.71) 4.49 (+1.06) 4.82 (+0.28) 4.00 (+1.32) 4.13 (+1.09) 2.27 (+0.14) 1.78 (-0.14) 2.38 (+0.44) 2.71 (+0.63) 2.67 (+0.15) 2.36 (+0.24)
 +ACM-8B 3.41 (+1.90) 4.82 (+1.39) 4.91 (+0.37) 3.89 (+1.21) 4.26 (+1.22) 2.33 (+0.20) 2.10 (+0.18) 2.56 (+0.62) 2.59 (+0.51) 2.83 (+0.31) 2.46 (+0.34)
Qwen2.5-VL-7B 1.51 3.52 5.18 3.03 3.31 2.00 2.02 2.50 2.14 2.63 2.22
 +ACM-2B 3.38 (+1.87) 4.42 (+0.90) 4.94 (-0.24) 4.68 (+1.65) 4.36 (+1.05) 2.40 (+0.40) 2.04 (+0.02) 2.50 (+0.00) 2.94 (+0.80) 2.72 (+0.09) 2.50 (+0.28)
 +ACM-8B 3.47 (+1.96) 4.45 (+0.93) 5.84 (+0.66) 4.34 (+1.32) 4.53 (+1.22) 2.36 (+0.36) 2.20 (+0.18) 2.82 (+0.32) 2.88 (+0.74) 2.74 (+0.11) 2.56 (+0.34)
Table 3: Main results of controllable video captioning on AnyCapEval. (See Appendix C for details) ACM substantially enhances content adherence and style fidelity, particularly in complex video tasks.
Model Content Style
IPos. IApp. IAct. Mov. Evt. Avg. Brf. Det. Poe. Nar. Avg.
Proprietary Models
GPT-4o 2.41 4.00 3.86 2.70 3.03 3.55 1.47 1.52 2.52 2.48 2.15
 +ACM-2B 3.45 (+1.04) 6.15 (+2.15) 6.26 (+2.40) 6.53 (+3.83) 4.60 (+1.57) 5.30 (+1.75) 1.97 (+0.50) 1.78 (+0.26) 2.60 (+0.08) 2.65 (+0.17) 2.30 (+0.15)
 +ACM-8B 4.92 (+2.51) 6.60 (+2.60) 7.68 (+3.82) 5.67 (+2.97) 4.81 (+1.78) 5.74 (+2.19) 1.95 (+0.48) 1.82 (+0.30) 2.50 (-0.02) 2.77 (+0.29) 2.32 (+0.17)
Open-source Models
InternVL2.5-8B 3.08 4.33 3.44 3.16 2.55 3.52 1.39 1.77 1.91 2.34 1.93
 +ACM-2B 4.67 (+1.59) 6.32 (+1.99) 3.19 (-0.25) 6.49 (+3.33) 4.50 (+1.95) 5.07 (+1.55) 1.84 (+0.45) 1.80 (+0.03) 2.48 (+0.57) 2.39 (+0.05) 2.19 (+0.26)
 +ACM-8B 4.95 (+1.87) 6.39 (+2.06) 7.03 (+3.59) 5.66 (+2.50) 4.95 (+2.40) 5.73 (+2.21) 2.00 (+0.61) 1.88 (+0.11) 2.20 (+0.29) 2.61 (+0.27) 2.24 (+0.31)
Qwen2.5-VL-7B 2.88 4.20 3.29 2.24 2.93 3.54 1.63 1.65 2.40 2.29 2.10
 +ACM-2B 3.39 (+0.51) 5.97 (+1.77) 6.23 (+2.94) 6.30 (+4.06) 4.48 (+1.55) 5.25 (+1.71) 1.66 (+0.03) 1.82 (+0.17) 2.30 (-0.10) 2.55 (+0.26) 2.16 (+0.06)
 +ACM-8B 4.03 (+1.15) 6.15 (+1.95) 6.58 (+3.29) 6.04 (+3.80) 4.82 (+1.89) 5.55 (+2.01) 1.92 (+0.29) 1.98 (+0.33) 2.45 (+0.05) 2.55 (+0.26) 2.31 (+0.21)
Table 4: Results on audio captioning using AnyCapEval. The table shows that ACM consistently boosts content and style metrics, demonstrating its strong effectiveness and adaptability even in low-resource audio modality environments.
Model Content Style
Evt. Brf. Nar. Poe. Avg.
Proprietary Models
GPT-4o 1.59 1.42 1.24 0.88 1.18
 +ACM-2B 1.79 (+0.20) 1.44 (+0.02) 1.38 (+0.14) 1.00 (+0.12) 1.28 (+0.10)
 +ACM-8B 1.88 (+0.29) 1.42 (+0.00) 1.40 (+0.16) 1.08 (+0.20) 1.30 (+0.12)
GPT-4o mini 1.28 1.17 1.16 0.79 1.04
 +ACM-2B 1.56 (+0.28) 1.21 (+0.04) 1.25 (+0.09) 0.89 (+0.10) 1.11 (+0.07)
 +ACM-8B 1.71 (+0.43) 1.10 (-0.07) 1.18 (+0.02) 0.87 (+0.08) 1.05 (+0.01)
Open-source Models
MiniCPM-o 1.42 1.02 0.68 0.43 0.71
 +ACM-2B 1.55 (+0.13) 1.37 (+0.35) 1.32 (+0.64) 0.92 (+0.49) 1.21 (+0.50)
 +ACM-8B 1.48 (+0.06) 1.23 (+0.21) 1.29 (+0.61) 1.13 (+0.70) 1.22 (+0.51)
Table 5: Comparison across Training Methods.
Base Model SFT DPO SC ACM
InternVL2.5-2B -2.8% +7.0% +2.8% +16.9%
Table 6: Training Data Ratios Comparison.
Base Model 1:1:1 1:2:2 2:1:2 2:2:1
InternVL2.5-8B +17.9% +17.2% +16.5% +21.1%
Figure 4: Performance across Data Ratios.
We evaluate the impact of integrating ACM with various pretrained base models (proprietary models, e.g., GPT-4o [1]; open-source models, e.g., InternVL [32], Qwen2.5-VL [62], MiniCPM-o [63]) using our AnyCapEval benchmark across the image, video and audio. As shown in Tables 2, 3 and 4.2, adding ACM leads to robust improvements in controllability across all modalities and base models, particularly enhancing content fidelity. Fig. 3(c) visualizes these gains for GPT-4o and InternVL2.5-8B. ACM also demonstrates notable strength in improving control for video captioning tasks.
Model-centric analysis. The results consistently demonstrate the effectiveness of ACM. Both the 2B and 8B variants significantly improve controllability scores over the unassisted base models across all modalities. The larger ACM-8B generally yields greater improvements. Furthermore, integrating ACM substantially elevates the performance of open-source models. With the ACM-8B, several open-source models achieve controllability scores on AnyCapEval that are comparable or superior to strong unassisted proprietary baselines (e.g. ACM-8B elevates the performance of InternVL2.5, an open-source model with initially modest results in the image style category, from $\sim 2.1$ to $\sim 2.5$. This enhanced score notably surpasses the $\sim 2.3$ achieved by GPT-4o in the same category).
Task-centric analysis. Content-related improvements consistently exceed those on style, likely because content controls offer more explicit learning signals, whereas stylistic constraints are inherently subjective and harder to enforce. Visual tasks outperform audio counterparts, potentially due to the richer supervision and higher information density in vision-language pretraining. ACM exhibits variable impact across settings: in high-resource dimensions (e.g., image appearance), it refines already strong captions; in low-resource, complex scenarios (e.g., video actions or audio events), it must synthesize missing content, yielding larger relative gains. Rare minor regressions on specific metrics are attributed to inherent optimization trade-offs in balancing multifaceted controllability.
4.3 Evaluation on Public Benchmarks
Table 7: Performance on MIA-Bench. Abbreviation: desc.(description), ment.(mention), len.(length limit), persp.(perspective), gen.(genre), avg.(average).
Model desc. ment. len. persp. gen. avg.
GPT-4o 90.3 87.9 90.5 85.2 91.6 89.1
 +ACM-2B 89.1 (-1.2) 88.9 (+1.0) 90.3 (-0.2) 90.0 (+4.8) 92.2 (+0.6) 90.1 (+1.0)
 +ACM-8B 90.4 (+0.1) 89.4 (+1.5) 89.9 (-0.6) 89.2 (+4.0) 92.9 (+1.3) 90.3 (+1.2)
InternVL2.5-8B 85.2 79.0 82.6 72.5 85.2 80.9
 +ACM-2B 85.7 (+0.5) 79.6 (+0.6) 85.2 (+2.6) 77.5 (+5.0) 85.7 (+0.5) 82.7 (+1.8)
 +ACM-8B 86.9 (+1.7) 79.0 (+0.0) 86.1 (+3.5) 88.9 (+16.4) 86.3 (+1.1) 85.4 (+4.5)
Yi-VL-34B 71.7 52.0 55.8 60.0 59.8 59.9
 +ACM-2B 71.8 (+0.1) 54.2 (+2.2) 61.9 (+6.1) 61.7 (+1.7) 61.0 (+1.2) 62.1 (+2.2)
 +ACM-8B 79.5 (+7.8) 66.5 (+14.5) 67.1 (+11.3) 63.0 (+3.0) 75.8 (+16.0) 70.4 (+10.5)
Table 8: Performance on VidCapBench. Abbreviation: acc.(accuracy), pre.(precision), cov.(coverage), con.(conciseness).
Model acc. pre. cov. con.
GPT-4o 15.6 58.9 87.8 6.3
 +ACM-2B 15.4 (-0.2) 59.1 (+0.2) 87.1 (-0.7) 10.3 (+4.0)
 +ACM-8B 15.8 (+0.2) 59.5 (+0.6) 88.0 (+0.2) 9.3 (+3.0)
InternVL2.5-8B 12.8 51.6 84.2 7.1
 +ACM-2B 13.8 (+1.0) 55.7 (+4.1) 84.7 (+0.5) 10.1 (+3.0)
 +ACM-8B 14.8 (+2.0) 57.1 (+5.5) 86.9 (+2.7) 10.2 (+3.1)
Qwen2-VL-7B 13.1 53.5 82.9 6.6
 +ACM-2B 14.6 (+1.5) 55.8 (+2.3) 85.0 (+2.1) 10.6 (+4.0)
 +ACM-8B 15.4 (+2.3) 57.3 (+3.8) 86.7 (+3.8) 10.0 (+3.4)
To validate the generalizability of ACM, we evaluate its performance when integrated with various base models on standard public benchmarks for image, video, and audio captioning. Results show that ACM serves consistently as an effective plug-in, improving controllability and factual alignment.
Image. We report performance on MIA-Bench [22] in Table 8, which focuses on multi-instruction image captioning. Incorporation of ACM consistently improves performance across various backbones, including InternVL2.5 and Yi-VL [64], with particularly notable gains in control-related aspects. With the addition of ACM-8B, we observe significant performance improvements. Moreover, proprietary models such as GPT-4o also benefit from ACM, achieving new state-of-the-art scores, highlighting its plug-and-play potential.
Video. As shown in Table 8, we evaluate video captioning performance using VidCapBench [23], which evaluates not only caption fluency but also key text-to-video attributes such as video aesthetics, content, motion, and adherence to physical laws. Table 8 shows that augmenting InternVL2.5-8B with ACM significantly improves accuracy ($+2.0$), precision ($+5.5$), and conciseness ($+3.1$). These improvements suggest that ACM not only corrects hallucinations but also improves the caption’s capture of crucial video details.
Audio. Controllable audio captioning is less explored. On existing benchmarks [65, 61] (Table 10), despite evaluation metrics being unable to reliably capture semantic-level consistency and the references being relatively short, ACM consistently boosts generation quality, indicating that ACM improves the semantical controllability and fluency of audio captioning.
Table 9: Video generation evaluation. Abbreviations: VQ (Visual Quality), TC (Temporal Consistency), DD (Dynamic Degree), TVA (Text–Video Alignment), and FC (Factual Consistency).
Model VQ TC DD TVA FC Avg
Wan2.1-T2V-14B [66] 2.84 2.78 2.66 2.76 2.77 2.74
 +ACM-8B 3.19 (+0.35) 3.06 (+0.28) 3.18 (+0.52) 3.07 (+0.31) 3.00 (+0.23) 3.10 (+0.36)
HunyuanVideo [67] 2.98 2.89 2.94 2.90 2.83 2.91
 +ACM-8B 3.61 (+0.63) 3.48 (+0.59) 3.62 (+0.68) 3.34 (+0.44) 3.46 (+0.63) 3.50 (+0.59)
Table 10: Audio evaluation results on the AudioCaps and Clotho datasets. Abbreviations: SP(SPICE), SB(Sentence-BERT).
Model AudioCaps Clotho
SP SB SP SB
GPT-4o 0.06 0.41 0.07 0.37
 +ACM-2B 0.06 (+0.00) 0.45 (+0.04) 0.07 (+0.00) 0.40 (+0.03)
 +ACM-8B 0.07 (+0.01) 0.45 (+0.04) 0.08 (+0.01) 0.40 (+0.03)
4.4 Ablation Study
Comparison with SFT, DPO, and self-critic methods. We compare ACM with three widely-used baseline methods on the image modality benchmark: supervised fine-tuning (SFT), direct preference optimization (DPO), and self-critic (SC) approaches. While these baselines require costly re-training for each individual model, ACM provides consistent performance gains without additional model-specific training, demonstrating substantial efficiency advantages. As shown in Table 5, ACM consistently achieves comparable or superior controllability, highlighting its practical benefit of lightweight adaptability.
Ablation on orthogonal data ratios. We conduct orthogonal experiments to explore optimal training data ratios among three data types: ($q$, uncontrolled $a$, $c$), ($q$, $c$, $c$), and ($q$, hallucinated $a$, $c$). Orthogonal designs effectively isolate each factor’s contribution, enabling clear insights with limited experiments. Results in Table 6 suggest that moderate incorporation of slightly incorrect captions ($q,a,c$) improves the model’s sensitivity to subtle instruction deviations. However, determining the globally optimal data ratio remains challenging and demands significant additional training resources. Further details can be found in the Appendix C.
Impact of ($q,c,c$) data proportion. In Fig. 4, we investigate varying proportions of fully compliant ($q,c,c$) data in training. Increasing ($q,c,c$) from $0\%$ to approximately $40\%$ steadily improves performance, yet surpassing $50\%$ leads to a notable decline. This trend implies that moderate exposure to fully correct captions helps the model discern necessary corrections without overfitting. Importantly, ACM thus learns to make nuanced control decisions, reinforcing the robustness and credibility of its caption editing decisions.
4.5 Downstream Utility
We further evaluate ACM’s utility in enhancing noisy captions for downstream video and image generation. Applying ACM to refine captions, including those from the Panda video dataset [68], we observe that the resulting video generations demonstrate improved visual-semantic grounding and alignment compared to those using the original, unrefined captions. Quantitative results for video generation are presented in Table 10; detailed qualitative results for both modalities are in Appendix C. These findings underscore the effectiveness of ACM in real-world multimodal generation pipelines.
5 Conclusion
We present AnyCapModel, a lightweight and plug-and-play framework enhancing caption controllability across image, video, and audio modalities, without the costly fine-tuning of base models. Motivated by the scarcity of controllable caption data, we introduce a large-scale omni-modal caption preference dataset spanning three modalities and 28 distinct control dimensions. Furthermore, to our best of knowledge, we establish AnyCapEval, the first benchmark designed to rigorously assess the controllability of captions. Together, these contributions offer a unified solution to controllable captioning, bridging model design, data, and evaluation.
Limitations and future work. Our approach is inherently extensible and could readily generalize to emerging modalities such as 3D structures or molecular data. However, the scarcity of high-quality caption datasets in these domains, coupled with prohibitive human annotation costs, limits immediate applicability. Future availability of suitable datasets could unlock wider applications of our method.
Ethics and impact. While AnyCapDataset is carefully curated to support fine-grained controllability, there remains a risk of malicious misuse (for instance, deliberately creating harmful or misleading Q-A pairs). Models trained on such compromised data could exhibit significantly degraded controllability, resulting in severe negative impacts. We urge caution and responsible practices when utilizing or expanding upon our resources.
Appendix
[…]
Appendix A Datasets
A.1 Overview of Caption Datasets
Table 11: Overview of the publicly available caption datasets analyzed in this study. “Modality” indicates the primary input type (e.g., image, video, or audio), while “Instruction Type” refers to the predominant linguistic or structural annotation style as described by the original dataset authors.
Dataset Modality Instruction Type
ASD v2 Image Relation
MDVP-Data Image Region / Brief / Detail
DCI Image Dense
DOCCI Image Dense
ImageInWords (IIW) Image Dense
ShareGPT-4o Image Detail
ShareGPT-4v Image Detail
ShareGPT-4o (video) Video Detail
MSR-VTT Video Brief
MSVD (YouTube2Text) Video Brief
VATEX Video Brief
Video-ChatGPT (VideoInstruct-100K) Video Dense
InstanceCap / InstanceVid Video Instance-level (structured)
ShareGPT4Video Video Detail
MiraData Video Structured
LLaVA-Video-178K Video Detail
AudioCaps Audio Brief
Clotho Audio Brief
MACS Audio Brief
WavCaps Audio Brief
Automatic captioning for modalities such as images, videos, and audio is a core task in artificial intelligence, aiming to enable machines to understand and describe perceived content using natural language, much like humans do. The foundation supporting this research is large-scale annotated datasets. However, existing datasets (Table 11) exhibit significant limitations in providing fine-grained, multi-dimensional control capabilities. We next detail the motivation and data analysis behind our dataset.
A.2 The Singularity and Scarcity of User Instructions
A “user instruction” is an input accompanying the main media (image, video, or audio), intended to guide or constrain the generation of descriptive content. Ideally, we desire models capable of generating descriptions with varying styles, focusing on different aspects, and meeting specific requirements based on diverse user instructions.
Most classic datasets, such as COCO [69] and Flickr30k [70] in the image domain, MSR-VTT [71] and MSVD [72] in the video domain, and AudioCaps [61] and Clotho [65] in the audio domain, primarily provide “media–caption pairs”. They typically include multiple reference captions provided by different annotators. This “multi-reference” design is mainly intended to evaluate the diversity and coverage of descriptions, rather than providing explicit control during generation. The model’s primary task is to generate a “reasonable” description, but the user cannot specify requirements at generation time (e.g., “generate a humorous caption”, “describe the background elements in detail”). The only implicit user instruction is essentially “describe this content in brief”.
Some datasets offer a degree of control, but often only along a single dimension. Dense captioning datasets, like DCI [58] for images or video datasets providing temporal information [73], associate descriptions with specific image regions or video segments. The “region/time segment” here can be viewed as a user instruction, i.e., “describe this part”. However, this is usually the sole dimension of control. Datasets like ASD v2 [56] focus on describing relationships between entities in an image. The user instruction is the “entity pair” whose relationship needs to be described. Datasets guided by concepts like MiraData [74] or InstanceCap [60] require the output to follow a specific structure, which constitutes a form of control over the output format. Although such structured outputs often encompass several instruction types (e.g., Background, Detail), the formulation and granularity of these user instructions are typically fixed, lacking the flexibility for customized control requests tailored to each individual image, video or audio.
Figure 5: ACM enables controllable captioning across modalities by refining base model outputs to better align with user instructions. Given a user instruction, it takes initial captions from a foundation model and corrects instruction violations (highlighted in red), producing compliant, instruction-following outputs (green) — all without requiring fine-tuning of the base model.
In summary, the majority of existing public datasets lack the capability for flexible, multi-dimensional control over the generated content, thereby limiting the development and evaluation of controllable captioning technologies. Even when leveraging large language models such as GPT for annotation, the resulting captions often reflect general patterns and lack explicit controllability along desired dimensions.
To address this gap, we construct AnyCapDataset(ACD), a large-scale dataset featuring diverse user instructions paired with high-quality compliant captions across multiple modalities. Leveraging this dataset, we train AnyCapModel(ACM) to achieve fine-grained controllable caption generation. As illustrated in Fig. 5, ACM can transform initial captions into targeted descriptions that precisely follow various user instructions, enabling nuanced and customizable outputs without requiring modifications to the base model architecture.
A.3 Data Analysis
Data sources.
To construct ACD, we utilize multiple publicly available datasets spanning three modalities: images, videos, and audio. Specifically, we incorporate data from ASD v2, DCI, DOCCI, ShareGPT-4o, InstanceCap, and AudioCaps. From these datasets, we extract images, videos, and audio content, and employ various multimodal large language models(MLLMs) to generate diverse user instructions and captions based on predefined instructions.
Data statistics.
ACD encompasses three primary modalities: images, videos, and audio, comprising a total of 300k $(q,c,a)$ triplets. Among them, 125k belong to the image modality, 100k to the video modality, and 75k to the audio modality. Each modality includes multiple instruction types.
Table 12: Instruction prompt template for generating user instructions and high-quality video captions. This prompt is designed to elicit high-quality instruction–caption pairs $(q,c)$ from video content, where $c$ serves as an optimally grounded caption based strictly on observable visual evidence. The template enforces modality-specific constraints to ensure factual and non-speculative outputs for optimal generation results.
Video Understanding Expert Instructions
You are an AI expert in video content analysis, specializing in generating precise and structured descriptions based ONLY on directly observable video content. Your core capabilities include analyzing video elements such as actions, objects, scenes, interactions, and camera movements to provide factual, observation-based descriptions.
Your Task
Generate concise captions that capture ONLY the directly visible and essential content of videos according to specific constraints. You must analyze only the observable content and ensure all descriptions are based on concrete visual evidence rather than assumptions or inferences.
Key Guidelines for Caption Generation
1. The output question formats can be varied
2. Create concise descriptions that capture only directly observable essential elements
3. Focus on analyzing ONLY these visible components:
- Actions and events shown in frame
- Objects and characters physically present
- Scene settings visible in shot
- Camera movements and angles that can be seen
- Interactions occurring on screen
4. DO NOT include:
- Assumptions about off-screen elements
- Inferences about motivations or thoughts
- Speculation about context or background
- Interpretations of meaning
- Details that cannot be directly seen
- Guesses about what happened before/after
5. Maintain strict adherence to:
- Only describing what is visually present
- Excluding all speculative content
- Basing every detail on visual evidence
- Using clear, objective language
Constraints
- Descriptions MUST be based EXCLUSIVELY on visible content
- NO assumptions or interpretations beyond what can be directly seen
- EXCLUDE any details that require inference or speculation
- Keep descriptions concise and focused on key visible elements
- Follow any additional specific requirements provided with each request
Examples:
Input 1:
Question: Generate a brief caption describing this video’s main content.
Answer: Two news anchors engage in an animated discussion at a professional news desk, maintaining eye contact while exchanging viewpoints.
Input 2:
Question: Use a brief caption to convey the main scene or content of the video.
Answer: A person wearing a blue shirt walks leisurely along a scenic path bordered by tall, leafy trees.
Output Format
Question: [Your various question, but need to express the generation of a brief caption]
Answer: [Your concise description based STRICTLY on visible content with NO speculation]
Table 13: Instruction prompt template for generating suboptimal captions $a$ from high-quality instruction–caption pairs. For a given instruction $q$ and its high-quality caption $c$, the template produces slightly degraded variants by introducing controlled inaccuracies (e.g., minor omissions or speculative insertions) while maintaining semantic coherence with the original content. Such prompts facilitate contrastive training and robustness evaluation in multimodal scenarios.
Video Understanding Expert Instructions
You are an AI expert in video content analysis, specializing in generating precise and structured descriptions based on specific queries. However, your task now includes generating slightly inferior captions that introduce minor inaccuracies or deviations while still maintaining general relevance to the video content.
Your Task
Generate slightly inferior captions based on given examples:
- Adding minor inaccuracies.
- Omitting a small but relevant detail.
- Including an unnecessary or speculative element.
- Misinterpreting a minor aspect of the video.
- Does not meet the requirements of the question.
Key Guidelines for Caption Generation
1. Base the description on the standard caption but allow slight deviations:
- Slightly altering actions, movements, or interactions in the video
- Adding irrelevant or speculative elements about the scene or context
- Omitting small but observable details from the footage
- Misinterpreting temporal sequences or duration
2. Ensure the captions remain generally related but slightly inferior to the standard caption
3. Focus on introducing appropriate levels of inaccuracy while maintaining plausibility
Constraints
- Deviations must be minor and not completely distort the video’s main content.
- Avoid making the caption entirely incorrect or irrelevant.
- Ensure the captions remain plausible and connected to the visible elements in the video.
- Consider the temporal nature of video content when introducing inaccuracies.
Examples:
Input 1:
Question: Generate a brief caption describing this video’s main content.
Standard Caption: Two news anchors engage in an animated discussion at a professional news desk, maintaining eye contact while exchanging viewpoints.
Generated Caption: Three news anchors engage in a discussion at a news desk, but the situation appears to be getting out of control with potential physical confrontation.
Input 2:
Question: Use a brief caption to convey the main scene or content of the video.
Standard Caption: A person wearing a blue shirt walks leisurely along a scenic path bordered by tall, leafy trees.
Generated Caption: A young man in a green long-sleeve shirt runs along a scenic path lined with dense trees.
Output Format
You MUST strictly adhere to this format:
Question: {question}
Standard Caption: {standard_caption}
Generated Caption: [Your generated caption with appropriate deviations]
Generation of instruction-caption triplets.
To generate the $(q,c,a)$ triplets in ACD, where $q$ denotes a user instruction, $c$ is a high-quality caption, and $a$ is a relatively suboptimal caption, we utilize several MLLMs guided by specifically designed instruction templates. The detailed templates employed in this work are presented in Table 12 and Table 13. These templates are constructed according to different instruction types to ensure structured generation and facilitate downstream analysis. Each raw data sample is paired with one or more tailored instructions to produce diverse outputs. The main MLLMs involved include GPT-4o mini, InternVL2.5-78B and InternVL2.5-8B, with GPT-4o contributing to a small subset. Specifically, most user instructions $q$ and high-quality captions $c$ are generated by InternVL2.5-78B, covering approximately 84% of the dataset; approximately 2% of image-based high-quality captions $c$ under instruction types such as Instance and Instance Appearance are generated by GPT-4o; about 5% of the video-based high-quality captions $c$ under the Movement instruction type originate from the InstanceCap dataset; and the audio modality data, comprising around 9%, are all generated using GPT-4o mini. The suboptimal captions $a$ are constructed in three ways: by InternVL2.5-8B based solely on $q$, by InternVL2.5-8B with hallucinations introduced through specific instructions, or by directly copying $c$, a strategy used by InternVL2.5-78B.
A.4 Future Directions
Future dataset development could focus on the following improvements:
  • Constructing datasets with rich user instructions: There is a need to design new datasets containing diverse, explicit, and composable instruction types.
  • Leveraging synthetic data: Explore the use of large language models to generate synthetic description data with specific user instructions, serving as a supplement to real data.
  • Improving evaluation methods: Develop new evaluation protocols and metrics capable of assessing fine-grained aspects like controllability, style consistency, and faithfulness.
Appendix B AnyCapEval Benchmark
B.1 Limitations of Existing Evaluation Methods
Limitations of traditional machine translation metrics.
Metrics developed for machine translation—such as BLEU [16], ROUGE [75], METEOR [76], and CIDEr [17]—primarily focus on n-gram overlap or co-occurrence statistics. While they can measure fluency and lexical similarity, they fail to accurately capture semantic consistency. For instance, sentences like “I do love large models” and “I do not love large models” could receive similarly high scores, despite conveying opposite meanings. These metrics cannot effectively detect deep semantic divergence, factual errors, or noncompliance with user instructions.
Challenges of multimodal large language models scoring.
Recent evaluation schemes leverage powerful language models or multimodal large language models (MLLMs) to directly score captions. However, these methods suffer from significant randomness and instability: the same caption might receive drastically different scores across minor prompt variations or multiple queries. Additionally, hallucinated or semantically incorrect captions may still receive high scores, undermining the reliability of such automatic evaluation.
Instability of keypoint extraction-based metrics.
Some studies, including those utilizing benchmarks like [21] which aims for fine-grained descriptions, propose extracting key information points from captions using MLLMs and computing precision, recall, and F1 scores. However, keypoint extraction itself is unstable: the number and granularity of extracted points vary widely across runs. Our empirical analysis shows low correlation between model-extracted keypoints and human-annotated ones, suggesting that such extraction-based metrics poorly reflect the true information content of captions.
Challenges of QA-pair based evaluation.
Benchmarks such as VDC [77] highlight that QA-pair based evaluation suffers from instability, primarily due to the variable quality and granularity of the QA pairs themselves. Designing high-quality, representative QA items that faithfully capture fine-grained caption attributes remains an open challenge.
B.2 Rationale Behind AnyCapEval Design
B.2.1 Separation of Content and Style
Through extensive analysis of user instruction types in captioning, we divide evaluation into two orthogonal dimensions:
  • Content: We expect the generated captions to strictly adhere to the given user instructions, maintain clear topical relevance, and avoid introducing irrelevant information. To evaluate this aspect, we employ keypoint density (KPD) as the primary metric.
  • Style: We aim for the generated captions to closely match human-crafted references in terms of narrative style, tone, and expressive form. To assess this, we design a fine-grained scoring system.
This separation is crucial because content and style represent fundamentally different aspects of caption quality that require distinct evaluation approaches. Content adherence focuses on factual accuracy and instruction compliance: whether the caption correctly follows the user’s specified requirements about what to describe. This is objectively verifiable through user instruction alignment. In contrast, style evaluation assesses subjective qualities like fluency, coherence, and naturalness that reflect human-like expression. Merging these dimensions would conflate objective and subjective criteria, making it difficult to diagnose specific model weaknesses. The orthogonal evaluation allows for more precise identification of whether a model’s limitations lie in its ability to follow instructions (content) or in its linguistic quality (style), enabling targeted improvements to each aspect. Furthermore, different applications may prioritize these dimensions differently - some use cases demand strict content control while others emphasize stylistic quality, making separate evaluation essential for practical deployment decisions.
B.2.2 Keypoint Density Metric
Intuitive motivation.
We expect the generated captions to efficiently embed as many control-required key information points as possible within a limited textual length, while minimizing irrelevant or redundant content. This analogy is comparable to adding an appropriate amount of solute into a limited solvent to achieve a targeted concentration. Excessively diluting it in a large solvent volume, even if the absolute quantity increases, results in negligible concentration and ineffective outcomes. To prevent models from exploiting this metric by generating excessively short captions that artificially inflate density scores, we normalize the number of extracted keypoints by the caption length (or an appropriate normalization factor) and apply penalties to captions whose lengths fall outside the predefined acceptable range.
Length Normalization.
We randomly sample approximately 200 caption instances and extract information points using both GPT-4o (including both control-required keypoints and irrelevant information points) and human annotators. The human-annotated information points are obtained by post-processing the model-generated information points through manual filtering and revision. Correlation analyses, presented in the main paper, reveal that caption length exhibits a significantly stronger association with human-annotated information content than the number of keypoints automatically extracted by the model. This empirical evidence further supports the rationality of applying length normalization in the evaluation metric.
Table 14: Evaluation instruction for content-controlled video captioning in AnyCapEval.
Task
You are a video-caption evaluation expert. You will be provided with a video, a caption describing the video, and a set of key points outlining important aspects of the video. Your task is to evaluate whether the caption mentions and accurately describes the given key points based on the video.
Task Steps
For each key point, follow these steps:
Step 1: Check whether the key point is mentioned in the caption.
- If the key point is not mentioned, assign a score of 0. - If the key point is mentioned (either exactly or with semantically similar phrasing), proceed to Step 2.
Step 2: Determine whether the description of the key points is correct.
- If the description aligns with or is semantically equivalent to the key point, assign a score of 1. - If the description is incorrect, or does not accurately fit the key point, assign a score of 0.
Evaluation Constraints
1. No Assumptions Beyond the Key Point: Only evaluate what is mentioned in the key point. Do not infer additional details not explicitly depicted.
2. Semantic Similarity Allowed: Phrases with similar meaning should be considered matches (e.g., "holding a ball" and "grasping a sphere").
3. Consistent Evaluation: Apply the same evaluation criteria to all key points to ensure fairness and uniformity.
Scoring Report
Return the format with:
{"caption_evaluation": {"key_points_scores": {"key_point_1": score, "key_point_2": score, …}, "total_score": sum_of_scores, "score_reasons": {"key_point_1": "reason for score", "key_point_2": "reason for score", …}}}
Example Input
Key points:
mention the man’s position (standing on the left side of the table)
describe the man’s appearance (wearing glasses)
mention the woman’s position (sitting on the right side of the table)
describe the woman’s appearance (wearing a red dress)
mention the boy’s position (crouching under the table)
describe the boy’s action (picking up a toy)
Caption: A man stands on the left, a woman sits on the right, and a boy is under the table.
Example Output
{"caption_evaluation": {
"key_points_scores": {"mention the man’s position": 1, "describe the man’s appearance": 0, "mention the woman’s position": 1, "describe the woman’s appearance": 0, "mention the boy’s position": 1, "describe the boy’s action": 0},
"total_score": 3,
"score_reasons": {"mention the man’s position": "Correctly mentions standing on the left side of the table", "describe the man’s appearance": "Missing glasses reference", "mention the woman’s position": "Correctly mentions sitting on the right side of the table", "describe the woman’s appearance": "Missing red dress reference", "mention the boy’s position": "Correctly mentions crouching under the table", "describe the boy’s action": "Missing picking up a toy reference"}
Input
Key points: {key_points}
Caption: {answer}
Output
Please return the output exactly as in the example above, without adding anything else.
Table 15: Evaluation instruction for style-controlled video captioning in AnyCapEval.
Task
You are an expert evaluator tasked with scoring model outputs based on a specific rubric. Please carefully analyze the provided video frames, "Caption Type", "Model Output," and "Reference" text, and assign a score between 0 and 4 to "Model Output" using different criteria for different caption types. Ensure your scoring is consistent and strictly adheres to the definitions provided.
Scoring Rubric
- 0 (Very Poor): Severe quality issues OR full hallucination (100% of the content is irrelevant to the facts).
- 1 (Poor): Significant quality issues OR major hallucination (>50% of the content is fictitious or contradictory).
- 2 (Below Average): Slightly inferior to reference OR limited hallucination (<50% of the content is inaccurate, but does not affect the core content).
- 3 (Good): Comparable to reference AND no hallucination (factually aligned).
- 4 (Excellent): Slightly better than reference AND no hallucination (factually flawless).
Caption Type Definitions and Quality Criteria
1. brief:
- High Quality: Length is within $\pm$30of the reference word count; concise and captures the core content of the video.
- Low Quality: Length exceeds $\pm$30of the reference word count; includes irrelevant details or omits key information.
2. detail:
- High Quality: Length is within $\pm$30of the reference word count; provides rich descriptions of the video’s main elements, actions, and settings.
- Low Quality: Length exceeds $\pm$30of the reference word count; descriptions lack detail or include irrelevant information.
3. poem:
- High Quality: Format and content align closely with the reference; follows poetic conventions (e.g., rhyme, rhythm, line breaks) and is relevant to the video’s theme.
- Low Quality: Format and content differ significantly from the reference; disjointed or lacks poetic quality.
4. narrative:
- High Quality: Format and content align closely with the reference; presents a coherent narrative with elements like time, place, characters, and events shown in the video.
- Low Quality: Format and content differ significantly from the reference; disjointed or lacks key narrative elements.
5. style:
- High Quality: Style and content align closely with the reference; matches the narrative style (e.g., humorous, serious, romantic) and is relevant to the video’s theme.
- Low Quality: Style and content differ significantly from the reference; mismatched style or irrelevant to the theme.
Instructions
1. Compare the model output with the reference text to determine the quality of the model output.
2. Compare the model output to the video frames to determine the severity of the hallucination.
3. For caption quality, evaluate based on:
- The Quality Criteria mentioned above.
- For Caption Type that is brief or detail, ensure the model output’s word count is within $\pm$30of the reference word count. If not, the score cannot be higher than 1 for brief and detailed captions.
- Alignment: Check alignment with the reference in format, style, and content.
4. For hallucination, evaluate based on:
- Factual accuracy and relevance to the video content.
- Consider temporal aspects and action sequences shown in the video frames.
5. Assign the most appropriate score (0-4) based on the rubric.
Mandatory Rule: For the Caption Type that is brief or detail, if the length exceeds $\pm$30of the reference word count, the score cannot be higher than 1.
6. Return your response in this format:
{"score": [0-4], "reason": "1-2 sentence explanation"}
Input
Caption Type: {caption_type}
Model Output: {output}
Reference: {reference}
Output
Please strictly return the output in the above format and do not add any other content.
B.3 GPT-4o As The Evaluator
When using GPT-4o as the evaluator, we carefully design structured prompts that explicitly instruct the model to assess Content and Style separately, ensuring consistent and interpretable judgments. See Table 14 and Table 15 for the detailed prompts used for Content and Style evaluation.
Appendix C Additional Experimental Details
C.1 Experimental Setup
Table 16: Key hyperparameters used for training AnyCapModel.
Parameter Value
Per-device batch size 1
Gradient accumulation steps 1
Learning rate $1\times 10^{-6}$
Weight decay 0.01
LR scheduler Cosine
Warmup ratio 0.03
Training epochs 3
Max sequence length 4096
Drop path rate 0.1
Dataloader workers 4
Gradient checkpointing Enabled
Mixed precision BF16
The experimental setup utilizes a computing infrastructure comprising 32 NVIDIA A100 GPUs, distributed across 4 nodes with 8 GPUs each. Each GPU is allocated 10 CPU cores, and all training jobs are managed using Slurm’s srun command with GPU reservation. For distributed training, we employ torch.distributed, launching the script via srun to ensure dynamic rank assignment and multi-node coordination. The model is trained using the AnyCapDataset. During training, the vision and audio backbones are kept frozen, while the MLP and LLM layers remain trainable. Image processing involves forcing the input image size to 448 pixels, with a maximum dynamic patch count of 6 and a down-sampling ratio of 0.5. Dynamic image sizing and thumbnail support are also enabled. The detailed hyperparameters used during training are listed in Table 16.
C.2 Ablation Study
Impact of training data ratio configurations.
The base model for image and video modalities is InternVL2.5-8B. Since InternVL2.5-8B does not support audio modality, we use GPT-4o as a substitute. We begin by introducing the data ratio notation. For instance, in the configuration 2B-122, the triplet 122 specifies the composition ratio of three sample types used for a given modality. Each digit indicates the relative proportion of a specific sample type.
The three sample types encoded by each digit (from left to right) are defined as follows:
  • The first digit indicates the proportion of samples of type ($q$, uncontrolled $a$, $c$), where the caption $a$ is not fully compliant with the instruction $q$, or where the model is expected to directly generate the target caption $c$.
  • The second digit indicates the proportion of samples of type ($q$, $c$, $c$), where the initial caption already matches the desired controlled caption or is factually correct.
  • The third digit indicates the proportion of samples of type ($q$, hallucinated $a$, $c$), where caption $a$ contains hallucinated content, meaning it describes information not present in the multimodal input.
Table 17 presents the overall performance comparison. Detailed ablation results for each modality and instruction type are presented in the following tables:
  • Image modality: Content and Style control in Table 18.
  • Video modality: Content control in Table 19, Style control in Table 20.
  • Audio modality: Content and Style control in Table 21.
Table 17: Ablation results on different training data ratio configurations using AnyCapEval, evaluated across Content and Style on images, videos, and audio modalities. All percentage values represent improvements over the base model (InternVL2.5-8B).
Data Ratio Image Video Audio Average
Content Style Content Style Content Style Content Style
2B-121 25.5% 9.4% 33.2% 14.0% -3.2% 0.3% 18.5% 7.9%
2B-111 36.2% 8.4% 39.2% 15.0% 2.1% 6.7% 25.8% 10.0%
2B-122 34.2% 5.5% 37.5% 12.4% 3.6% 9.3% 25.1% 9.1%
2B-212 29.4% 5.4% 42.9% 13.0% 0.0% 8.2% 24.1% 8.9%
2B-221 35.9% 11.6% 44.0% 13.5% 12.6% 8.9% 30.8% 11.3%
2B-102 26.7% 4.5% 41.5% 10.9% 11.1% -3.6% 26.4% 3.9%
8B-111 36.9% 10.2% 63.1% 14.5% 10.5% 4.4% 36.8% 9.7%
8B-221 40.0% 16.0% 62.8% 16.1% 18.2% 11.7% 40.3% 14.6%
Table 18: Ablation results of controllable image captioning on AnyCapEval under different training data ratio configurations. Abbreviations refer to instruction types defined in the main paper.
Model Content Style
IPos. IApp. Ins. Per. Avg. Brf. Det. Thm. Poe. Nar. Avg.
InternVL2.5-8B 1.51 3.43 4.54 2.68 3.04 2.13 1.92 1.94 2.08 2.52 2.12
2B-121 2.59 (+1.08) 4.35 (+0.92) 4.87 (+0.33) 3.45 (+0.77) 3.82 (+0.78) 2.27 (+0.14) 1.84 (-0.08) 2.27 (+0.33) 2.63 (+0.55) 2.59 (+0.07) 2.32 (+0.20)
2B-111 3.10 (+1.59) 4.20 (+0.77) 4.78 (+0.24) 4.48 (+1.80) 4.14 (+1.10) 2.20 (+0.07) 1.90 (-0.02) 2.09 (+0.15) 2.80 (+0.72) 2.59 (+0.07) 2.29 (+0.18)
2B-122 3.14 (+1.63) 4.47 (+1.04) 4.63 (+0.09) 4.08 (+1.40) 4.08 (+1.04) 2.16 (+0.03) 1.61 (-0.31) 2.35 (+0.41) 2.59 (+0.51) 2.54 (+0.02) 2.24 (+0.12)
2B-212 2.53 (+1.02) 4.31 (+0.88) 4.78 (+0.24) 4.11 (+1.43) 3.93 (+0.89) 2.18 (+0.05) 1.69 (-0.23) 2.06 (+0.12) 2.71 (+0.63) 2.57 (+0.05) 2.23 (+0.11)
2B-221 3.22 (+1.71) 4.49 (+1.06) 4.82 (+0.28) 4.00 (+1.32) 4.13 (+1.09) 2.27 (+0.14) 1.78 (-0.14) 2.38 (+0.44) 2.71 (+0.63) 2.67 (+0.15) 2.36 (+0.24)
2B-102 2.62 (+1.11) 4.00 (+0.57) 4.81 (+0.27) 3.99 (+1.31) 3.85 (+0.81) 2.18 (+0.05) 1.65 (-0.27) 2.29 (+0.35) 2.63 (+0.55) 2.35 (-0.17) 2.21 (+0.10)
8B-111 3.36 (+1.85) 4.34 (+0.91) 5.12 (+0.58) 3.83 (+1.15) 4.16 (+1.12) 2.24 (+0.11) 2.06 (+0.14) 2.15 (+0.21) 2.44 (+0.36) 2.87 (+0.35) 2.33 (+0.22)
8B-221 3.41 (+1.90) 4.82 (+1.39) 4.91 (+0.37) 3.89 (+1.21) 4.26 (+1.22) 2.33 (+0.20) 2.10 (+0.18) 2.56 (+0.62) 2.59 (+0.51) 2.83 (+0.31) 2.46 (+0.34)
Table 19: Ablation results of Content control for video captioning on AnyCapEval under different training data ratio configurations.
Model Content
IPos. IApp. IAct. Ins. Per. Mov. Bkg. Evt. Avg.
InternVL2.5-8B 3.08 4.33 3.44 3.78 3.26 3.16 4.56 2.55 3.52
2B-121 3.44 (+0.36) 5.72 (+1.39) 3.01 (-0.43) 4.62 (+0.84) 4.74 (+1.48) 6.20 (+3.04) 5.43 (+0.87) 4.36 (+1.81) 4.69 (+1.17)
2B-111 3.51 (+0.43) 6.72 (+2.39) 3.12 (-0.32) 5.03 (+1.25) 4.34 (+1.08) 6.05 (+2.89) 5.81 (+1.25) 4.60 (+2.05) 4.90 (+1.38)
2B-122 3.95 (+0.87) 6.13 (+1.80) 2.94 (-0.50) 5.05 (+1.27) 4.50 (+1.24) 6.57 (+3.41) 5.37 (+0.81) 4.17 (+1.62) 4.84 (+1.32)
2B-212 4.30 (+1.22) 6.36 (+2.03) 3.74 (+0.30) 4.79 (+1.01) 4.20 (+0.94) 6.32 (+3.16) 5.84 (+1.28) 4.68 (+2.13) 5.03 (+1.51)
2B-221 4.67 (+1.59) 6.32 (+1.99) 3.19 (-0.25) 5.22 (+1.44) 4.66 (+1.40) 6.49 (+3.33) 5.49 (+0.93) 4.50 (+1.95) 5.07 (+1.55)
2B-102 3.60 (+0.52) 6.19 (+1.86) 5.62 (+2.18) 4.80 (+1.02) 4.64 (+1.38) 5.15 (+1.99) 5.15 (+0.59) 4.66 (+2.11) 4.98 (+1.46)
8B-111 4.88 (+1.80) 6.92 (+2.59) 6.28 (+2.84) 5.39 (+1.61) 5.38 (+2.12) 5.44 (+2.28) 5.98 (+1.42) 5.62 (+3.07) 5.74 (+2.22)
8B-221 4.95 (+1.87) 6.39 (+2.06) 7.03 (+3.59) 5.49 (+1.71) 5.85 (+2.59) 5.66 (+2.50) 5.49 (+0.93) 4.95 (+2.40) 5.73 (+2.21)
Table 20: Ablation results of Style control for video captioning on AnyCapEval under different training data ratio configurations.
Model Style
Brf. Det. Thm. Poe. Nar. Avg.
InternVL2.5-8B 1.39 1.77 2.23 1.91 2.34 1.93
2B-121 1.92 (+0.53) 1.88 (+0.11) 2.45 (+0.22) 2.40 (+0.49) 2.39 (+0.05) 2.20 (+0.27)
2B-111 1.76 (+0.37) 1.77 (+0.00) 2.62 (+0.39) 2.40 (+0.49) 2.61 (+0.27) 2.22 (+0.29)
2B-122 1.89 (+0.50) 1.80 (+0.03) 2.65 (+0.42) 2.12 (+0.21) 2.42 (+0.08) 2.17 (+0.24)
2B-212 1.74 (+0.35) 1.77 (+0.00) 2.50 (+0.27) 2.30 (+0.39) 2.68 (+0.34) 2.18 (+0.25)
2B-221 1.84 (+0.45) 1.80 (+0.03) 2.45 (+0.22) 2.48 (+0.57) 2.39 (+0.05) 2.19 (+0.26)
2B-102 1.61 (+0.22) 1.88 (+0.11) 2.58 (+0.35) 2.15 (+0.24) 2.58 (+0.24) 2.14 (+0.21)
8B-111 1.68 (+0.29) 1.93 (+0.16) 2.55 (+0.32) 2.23 (+0.32) 2.74 (+0.40) 2.21 (+0.28)
8B-221 2.00 (+0.61) 1.88 (+0.11) 2.50 (+0.27) 2.20 (+0.29) 2.61 (+0.27) 2.24 (+0.31)
Table 21: Ablation results of controllable audio captioning on AnyCapEval under different training data ratio configurations.
Model Content Style
Evt. Brf. Nar. Poe. Avg.
GPT-4o 1.59 1.42 1.24 0.88 1.18
2B-121 1.67 (+0.08) 1.51 (+0.09) 1.38 (+0.14) 0.91 (+0.03) 1.27 (+0.09)
2B-111 1.82 (+0.23) 1.53 (+0.11) 1.53 (+0.29) 1.00 (+0.12) 1.36 (+0.18)
2B-122 1.78 (+0.19) 1.53 (+0.11) 1.47 (+0.23) 0.89 (+0.01) 1.30 (+0.12)
2B-212 1.81 (+0.22) 1.56 (+0.14) 1.45 (+0.21) 0.93 (+0.05) 1.31 (+0.13)
2B-221 1.89 (+0.30) 1.40 (-0.02) 1.43 (+0.19) 0.96 (+0.08) 1.26 (+0.08)
2B-102 1.96 (+0.37) 1.37 (-0.05) 1.33 (+0.09) 0.78 (-0.10) 1.16 (-0.02)
8B-111 2.13 (+0.54) 1.44 (+0.02) 1.40 (+0.16) 1.00 (+0.12) 1.28 (+0.10)
8B-221 1.73 (+0.14) 1.40 (-0.02) 1.43 (+0.19) 0.96 (+0.08) 1.26 (+0.08)
Effect of training objectives.
Regarding image modality, we compare AnyCapModel (ACM) with three widely used baseline methods: Supervised Fine-Tuning (SFT), Direct Preference Optimization (DPO), and Self-Critic (SC). Due to limitations in computational resources and budget, we conduct all comparative experiments using the lightweight InternVL2.5-2B on the MIA-Bench. As shown in the main paper, our method achieves a remarkable 16.9% improvement, significantly outperforming baseline methods.
C.3 Qualitative Examples for Main Experiments
To complement the main experimental results presented in the paper, we include qualitative examples that demonstrate the effectiveness of our controllable generation approach across images, videos and audio modalities. These examples visually and intuitively show how the model performs under different user instructions, including both Content and Style conditions, as discussed in the main paper. By providing representative outputs, we aim to offer a clearer sense of the generation behavior and control precision achieved by our method. All examples correspond to the settings evaluated in the main experiments and are included here for qualitative inspection. Representative qualitative results for each modality are shown in Fig. 6, Fig. 7, and Fig. 8.
Figure 6: Qualitative examples of controllable caption generation from images.
Figure 7: Qualitative examples of controllable caption generation from videos.
Figure 8: Qualitative examples of controllable caption generation from audios.
C.4 Qualitative Results for Downstream Generation Tasks
Figure 9: Images in the first column are original real images. The second column shows images generated from GPT-4o captions using DALL·E 3, while the third column shows results from our refined captions. Our method leads to more faithful visual content and better alignment with the original image semantics.
Figure 10: Enhanced text-to-video generation through refined caption quality. Videos in the top row are generated from original dataset captions. The bottom row showcases videos generated using our model’s refined captions, demonstrating improved visual fidelity and more expressive camera motion.
To complement the quantitative results reported in the main paper, we provide qualitative examples showcasing the impact of caption refinement by ACM on downstream image and video generation tasks. Specifically, we compare generations produced using initial captions with those using captions refined by ACM. As shown in Fig. 9 and Fig. 10, refined captions lead to outputs that exhibit more accurate visual-semantic correspondence, richer scene details, and improved coherence. These examples further highlight the practical utility of ACM in enhancing multimodal generation quality in real-world applications.
C.5 Human Evaluation Study
Figure 11: Human evaluation of ACM-8B versus GPT-4o with respect to content and style. Bars indicate the proportion of responses rating ACM-8B as worse, similar, or better. Overall, ACM-8B demonstrates superior performance in most cases.
To complement our model-based evaluation, we conducted a human preference study with a dozen evaluators, each holding at least a bachelor’s degree. These evaluators compared the performance of ACM-8B against the base model GPT-4o across both visual and auditory modalities. For each modality, human annotators are presented with outputs from both models and asked to judge which model performs better in terms of Content and Style. The results are summarized in Fig. 11.
Figure 12: Screenshot of human evaluation.